A voltaic cell is constructed with two silver-silver chloride electrodes, each of which is based on the following half-reaction: AgCI(s) teAg(s) + CI (aq) The two cell compartments have [CI] = 0.0 150 mol L-i and [Cl]-2.55 mol Li, respectively. 1. Which electrode is the cathode of the cell? 2. What is the standard EMF of the cell? 3. What is the cell EMF for the concentrations given? 4. For each electrode, predict if [CI] will increase, decrease, or stay the same as the cell operates.
Answer #1

Answer:

1)

In this case, the AgCl electrode is only cathode

2)

the standard EMF of this cell is 0

as Ag + Cl-(anode) + AgCl ==> Ag + Cl-(cathode) + AgCl . . .Eo cell = +0.00 V

3) E cell = Eo cell - 0.059/n log Q = 0.00 V - 0.059/1 log ([Cl- cathode] / [Cl- anode])
= 0.00 V - 0.0295 log (3.00 / 0.0160) = 0.00 V - (0.0295) log (0.0150 / 2.55) = +0.0657 V

4)

Cl- concentration will increase on cathode and it will decrease on the anode

please rate it if it helps you

Thanks :)